Given that s = 1.59t(1 - 3v), obtain the value of v when s = 3.52 and t = 21.56
Answer by adamchapman(301)

s = 1.59t(1 - 3v)
plug in your values for s and t:
3.52=(1.59*21.56)(1-3v)
simplify and rearrange to find v:
3.52=(34.2804)(1-3v)
3.52/34.2804=1-3v
0.10268=1-3v
3v=1-0.10268
3v=0.8973
v=0.8973/3
v=0.2991
I hope this helps.
P.S. I am trying to start up my own homework help website. I would be extremely grateful if you would e-mail me some feedback on the help you received to adam.chapman@student.manchester.ac.uk
Adam